# Sequential Randomization Tests Using E-values: A Betting Approach for Clinical Trials

### v2

@@ Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We present a nonparametric sequential test – the randomization e-process (e-RT) – that derives validity solely from the randomization mechanism. Using a betting framework, e-RT constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. The e-RT provides a conservative, assumption-free complement to model-based sequential analyses.
@@ 1 Introduction
Sequential monitoring of randomized controlled trials requires methods that control Type I errors regardless of when or why the monitoring stops. Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries. When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guarantees may erode.
@@
We present the randomization e-process (e-RT), a sequential test that derives validity solely from the trial’s randomization mechanism. The construction uses a betting framework: at each enrollment, we wager on the treatment assignment given the observed outcome. Under the null hypothesis of no treatment effect, treatment assignment is independent of outcome, and no betting strategy can systematically grow wealth. This independence, guaranteed by randomization, is the only assumption required.
The contributions of this note are as follows.
  1. 1.
    An explicit construction of e-RT with proof of validity
  2. 2.
    Simulation studies demonstrating Type I error calibration and power
2 The e-RT Construction
@@ 2.1 Setup
Consider a sequential randomized trial. At each enrollment $i=1,2,\ldots$, we observe $T_{i}$: assignment of treatment (control or intervention) and $Y_{i}$: binary outcome (event or no event). Treatment is assigned with known probability $p$, typically $p=0.5$ for 1:1 randomization.
The null hypothesis is that treatment has no effect. Under this hypothesis, knowing the outcome of a patient tells you nothing about which arm they were assigned to – outcome and treatment are independent ($Y_{i}\perp T_{i}$).
2.2 Betting Construction
e-RT constructs a wealth process $W_{n}$ by betting sequentially on treatment assignments.
For each patient, after observing their outcome but before learning their treatment assignment, we place a bet. The bet is a fraction $\lambda_{i}\in[0,1]$ representing how much we wager on the patient being in the treatment arm.
Under the null hypothesis, outcome tells us nothing about assignment; hence we can’t systematically win. Under the alternative, outcomes differ by arm, so we can bet profitably.
@@ 2.2 Betting Construction
$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=\text{intervention}\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=\text{control}\end{cases}$ (1)
starting from $W_{0}=1$. Note that the intervention and control here are simply labels, they could be called whatever (A versus B, etc).
2.3 Adaptive Betting Strategy
The betting strategy uses accumulated data to inform each bet. Here is how it works:
  1. 1.
    A patient is randomized (assignment blinded).
  2. 2.
    Outcome is observed.
  3. 3.
    Historical records are checked. The most likely arm for this outcome is defined.
  4. 4.
    Bet is placed according to previous data.
  5. 5.
    Assignment is revealed.
  6. 6.
    Wealth updated. If guess is correct, it grows; if wrong, it shrinks.
Under the null hypothesis, any historical imbalance between arms is noise. Bets will be sometimes right, sometimes wrong, and wealth fluctuates randomly around 1. Under the alternative, the imbalance is real signal. Bets will be systematically correct, and wealth grows over time.
Formally, let:
$\hat{\delta}_{i-1}=\text{(event rate in treatment)}-\text{(event rate in control)}$ (2)
estimated from all patients before $i$. The betting fraction is:
$\lambda_{i}=\begin{cases}0.5+0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{event}\\ 0.5-0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{no event}\end{cases}$ (3)
where $c_{i}\in[0,1]$ is a confidence parameter that ramps up from 0 to 1 over a burn-in period, preventing wild bets when data are sparse:
$c_{i}=\min\left(1,\frac{i-\text{burn-in}}{\text{ramp}}\right)$ (4)
The above can be read as: At any specific point, the betting fraction $\lambda$ is defined as $0.5$ plus the wager if there has been an event and $0.5$ minus the wager if there hasn’t been an event. The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{i}$. This factor is simply a factor that avoids us from making full bets early on. It is determined by a burn-in period (where we don’t bet, just collect data), and a ramp, which is also arbitrarily defined, so that wager is "muffled" early on. Finally, the $0.5$ multiplier before $c_{i}$ and the wager ($\hat{\delta}_{i-1}$) simply keeps $\lambda$ between zero and one.
@@ 2.4 Worked Example
Consider a trial comparing intervention versus control, with a binary outcome (event or no event). Event is mortality, which is expected to be lower with intervention. Allocation is 1:1 ($p=0.5$). Assume burn-in is complete ($c=1$).
@@
Patient 200 has an event (dies). Where is this patient likely from? Events are more common in control (40.4% vs 35.0%), so probably control. We bet $\lambda=0.5+0.5\times(-0.054)=0.473$ on intervention and $1-\lambda=0.527$ on control. Assignment revealed and patient indeed is in control group. We guessed right! Our multiplier will be $0.527/0.5=1.054$, hence wealth grows 5.4%.
Now patient 201 has been enrolled. First, we update the counts with patient 200’s data. Intervention arm now has 100 patients, 35 events (rate = 35.0%), and control arm has 100 patients, 41 events (rate = 41.0%). The $\hat{\delta}_{200}=0.350-0.410=-0.060$. Patient 201 does not have an event. Non-events are more common in intervention (65.0% vs 59.0%), so probably intervention. We will bet once again. This time we bet $\lambda=0.5-0.5\times(-0.060)=0.530$ on intervention and $1-\lambda=0.470$ on control. Finally, assignment is revealed and we guessed right again: patient is in intervention. Our multiplier =0.530/0.5=1.060=0.530/0.5=1.060. Wealth grows another 6.0%.
Finally, patient 202 has been enrolled. We update counts once more with intervention having 101 patients and 35 events (rate = 34.7%) and control having 100 patients with 41 events (rate = 41.0%). Our $\hat{\delta}_{201}=0.347-0.410=-0.063$. Patient 202 has an event (dies). Events cluster in control, so we bet toward control: $\lambda=0.5+0.5\times(-0.063)=0.469$ on intervention and $1-\lambda=0.531$ on control. Assignment is revealed and we guessed wrong—patient is in intervention group! Our multiplier will be $0.469/0.5=0.938$. Wealth shrinks 6.2%.
Cumulative wealth is the product of all multipliers:
@@
Despite one wrong guess, wealth grew 4.8% over these three patients.
Even when intervention works, some intervention patients die. We guess wrong sometimes. But if intervention truly reduces mortality, events cluster in control and non-events cluster in intervention — we are right more often than wrong, and wealth grows on average.
Under the null hypothesis (no effect), right and wrong guesses balance out. Wealth fluctuates randomly around 1.
In summary: the outcome determines $\lambda$ (which direction to bet); the treatment arm determines which multiplier formula to apply (whether the bet paid off). The machinery is agnostic to whether the effect size is following the expected direction; it simply keeps betting according to what is more likely.
@@ 2.5 Validity
Under the null hypothesis, the wealth process $(W_{n})$ is a test martingale — its expected value cannot grow.
@@
Under the null, outcome and treatment are independent. So after seeing the outcome, treatment assignment is still just a coin flip with probability $p$. The expected wealth multiplier is:
$\mathbb{E}[\text{multiplier}]$ $=p\times\frac{\lambda}{p}+(1-p)\times\frac{1-\lambda}{1-p}$ (6)
$=\lambda+(1-\lambda)$ (7)
$=1$ (8)
No matter how we choose our bet $\lambda$, we can’t systematically grow wealth when the null is true. ∎
@@
For any stopping time $\tau$, $\mathbb{E}[W_{\tau}]\leq 1$ under the null. By Markov’s inequality, $P(W_{\tau}\geq 1/\alpha)\leq\alpha$. Thus rejecting when $W_{\tau}\geq 1/\alpha$ controls Type I error at level $\alpha$, regardless of the stopping rule.
@@ 3 Simulation Studies
We evaluated e-RT operating characteristics by simulation. For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size. We used burn-in = 50 patients and ramp = 100 patients. Control arm event rate was 40% in all scenarios.
@@ 3.1 Results
Table 1: e-RT Operating Characteristics
ARR Target Power $n$ Type I Error e-RT Power Median Crossing
@@
Power was approximately 50% for trials designed with 80% power, and 63–66% for trials designed with 90% power. When the e-RT rejected the null, it did so at approximately half the planned sample size (median crossing 47–56% of total enrollment).
@@ 3.2 Interpretation
The e-RT is not a replacement for the primary analysis—it is a continuous monitoring tool. A trial designed for 80% power with a traditional test gains a 50% chance of stopping early via e-RT, at roughly the halfway point. If the threshold is not crossed, the trial proceeds to completion and the planned analysis is conducted.
This represents a free option: anytime-valid monitoring with no alpha spending and no pre-specified interim looks. The “cost” is that the e-RT alone has lower power than a fixed-sample test. But when layered on top of a properly powered trial, it provides early stopping when effects are larger than anticipated.
@@ 3.3 Trajectory Examples
Figure 1 shows example e-RT trajectories from 30 simulated trials under the null hypothesis (both arms 40% event rate). Sample sizes correspond to trials designed for 80% power (n = 712, left) and 90% power (n = 954, right) to detect a 10% ARR. Under the null, wealth fluctuates randomly around 1. Some trajectories temporarily rise toward the threshold but eventually drift back down. The downward drift over time reflects the accumulating “cost” of betting on noise.
Figure 1: e-RT trajectories under the null hypothesis. Left: n = 712 (80% power design). Right: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Dotted gray line: neutral (wealth = 1). Under the null, no trajectory crosses the threshold.

### v3

# Sequential Randomization Tests Using e-values: Applications for trial monitoring

@@ Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We discuss a nonparametric sequential test and its application to continuous and time-to-event endpoints that derives validity solely from the randomization mechanism. Using a betting framework, these tests constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. These methods provide a conservative, assumption-free complement to model-based sequential analyses.
@@ 1 Introduction
Sequential monitoring of randomized controlled trials requires methods that control Type I errors regardless of when or why the monitoring stops. Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries. When these assumptions fail, or when trials adapt in ways not fully prespecified, validity guarantees may erode.
Monitoring clinical trials, specially in the context of acutely ill patients, is of paramount importance. Traditional methods include $\alpha-spending$ functions. When interim analyses are spaced, evidence grows unnoticed between the interim analyses. This may both delay the implementation of potentially useful strategies or prolong trials when safety signals arise early.
@@
pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which tests treatment effects by wagering on treatment assignments given observed outcomes. The intuition is discussed by ramdas2021game. Under the null hypothesis, randomization ensures that assignments are independent of outcomes, so no betting strategy can systematically accumulate wealth.
We discuss potential methods for trial monitoring based on previous efforts. For binary outcomes, our construction implements the i-bet framework with a specific adaptive betting strategy. We present methods with unified notation, validity proofs, and simulation studies demonstrating Type I error control and power characteristics.
2 Overall construction
@@ 2.1 Setup
Consider a sequential randomized trial with 1:1 allocation. At each enrollment $i=1,2,\ldots$, we observe:
  • $T_{i}\in\{0,1\}$: treatment assignment (0 = control, 1 = intervention)
  • $Y_{i}\in\{0,1\}$: binary outcome (0 = no event, 1 = event)
Treatment is assigned with known probability $p=P(T_{i}=1)$, typically $p=0.5$.
The null hypothesis is that treatment assignment has no effect on outcome:
$H_{0}:Y_{i}\perp T_{i}\text{ for all }i$ (1)
Under this hypothesis, observing the outcome provides no information about which arm the patient was assigned to.
2.2 Wealth Process
Following pmlr-v177-duan22a, we construct a wealth process by wagering on treatment assignments. After observing outcome $Y_{i}$ but before learning treatment assignment $T_{i}$, we choose $\lambda_{i}\in[0,1]$: the fraction wagered on intervention.
@@ 2.2 Wealth Process
$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=1\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=0\end{cases}$ (2)
starting from $W_{0}=1$. When we bet toward the correct arm, wealth grows; when wrong, it shrinks.
2.3 Betting Strategy
The validity guarantee holds for any betting strategy where $\lambda_{i}$ depends only on $\mathcal{F}_{i-1}$. Power depends on choosing bets that grow wealth under the alternative. We use a strategy that learns the treatment effect from accumulating data.
Let:
$\hat{\delta}_{i-1}=\text{(event rate in intervention)}-\text{(event rate in control)}$ (3)
estimated from patients $1,\ldots,i-1$. The betting fraction is:
$\lambda_{i}=\begin{cases}0.5+0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=1\\ 0.5-0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=0\end{cases}$ (4)
where $c_{i}\in[0,1]$ ramps from 0 to 1 over a burn-in period:
$c_{i}=\min\left(1,\max\left(0,\frac{i-n_{0}}{n_{r}}\right)\right)$ (5)
with $n_{0}$ the burn-in period and $n_{r}$ the ramp period. This prevents large bets when $\hat{\delta}_{i-1}$ is unstable due to small samples.
The logic: if $\hat{\delta}>0$ (more events in intervention), then events suggest intervention and non-events suggest control. If $\hat{\delta}<0$ (fewer events in intervention), then events suggest control and non-events suggest intervention. The factor of $0.5$ before $c_{i}\cdot\hat{\delta}_{i-1}$ ensures $\lambda_{i}\in[0,1]$.
@@ 2.4 Worked Example
Consider a trial comparing intervention versus control, with a binary outcome (event or no event). Event is mortality, which is expected to be lower with intervention. Allocation is 1:1 ($p=0.5$). Assume burn-in is complete ($c_{i}=1$).
@@
Patient 200 has an event (dies). Where is this patient likely from? Events are more common in control (40.4% vs 35.0%), so probably control. We bet $\lambda=0.5+0.5\times(-0.054)=0.473$ on intervention and $1-\lambda=0.527$ on control. Assignment revealed: control. We guessed right. Multiplier: $0.527/0.5=1.054$. Wealth grows 5.4%.
Patient 201: Update counts—intervention has 100 patients, 35 events (35.0%); control has 100 patients, 41 events (41.0%). $\hat{\delta}_{200}=-0.060$. Patient 201 has no event. Non-events are more common in intervention (65.0% vs 59.0%), so probably intervention. Bet: $\lambda=0.5-0.5\times(-0.060)=0.530$. Assignment revealed: intervention. Multiplier: $0.530/0.5=1.060$. Wealth grows 6.0%.
Patient 202: Update counts—intervention has 101 patients, 35 events (34.7%); control has 100 patients, 41 events (41.0%). $\hat{\delta}_{201}=-0.063$. Patient 202 has an event. Bet toward control: $\lambda=0.5+0.5\times(-0.063)=0.469$. Assignment revealed: intervention. Wrong guess. Multiplier: $0.469/0.5=0.938$. Wealth shrinks 6.2%.
Cumulative wealth:
@@
Despite one wrong guess, wealth grew 4.8% over these three patients. Under the alternative, correct guesses outnumber incorrect ones on average and wealth grows. Under the null, right and wrong guesses balance out and wealth fluctuates around 1.
@@ 2.5 Validity
Under the null hypothesis, the wealth process $(W_{n})$ is a nonnegative martingale.
@@
Under the null, outcome and treatment are independent. After observing outcome $Y_{i}$, treatment assignment remains a coin flip with $P(T_{i}=1)=p$. The expected multiplier given any bet $\lambda_{i}$ is:
$\mathbb{E}[\text{multiplier}\mid\lambda_{i}]$ $=p\times\frac{\lambda_{i}}{p}+(1-p)\times\frac{1-\lambda_{i}}{1-p}$ (7)
$=\lambda_{i}+(1-\lambda_{i})=1$ (8)
Thus $\mathbb{E}[W_{i}\mid W_{i-1},\lambda_{i}]=W_{i-1}$. ∎
@@
By Ville’s inequality (ville1939), for any nonnegative martingale starting at 1:
$\Pr_{H_{0}}\left(\sup_{n\geq 1}W_{n}\geq\frac{1}{\alpha}\right)\leq\alpha$ (9)
Thus rejecting when wealth ever exceeds $1/\alpha$ controls Type I error at level $\alpha$, regardless of when or why monitoring stops.
@@ 3 Simulation Studies
We evaluated operating characteristics by simulation. For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size. We used burn-in = 50 patients and ramp = 100 patients. Control arm event rate was 40% in all scenarios.
@@ 3.1 Results
Table 1: Operating Characteristics for Binary e-process
ARR Target Power $n$ Type I Error Binary e-process Power Median Crossing
@@
Power was approximately 50% for trials designed with 80% power, and 63–66% for trials designed with 90% power. When the process rejected the null, it did so at approximately half the planned sample size (median crossing 47–56% of total enrollment).
@@ 3.2 Interpretation
This method seems to be appealing as a continuous monitoring tool. A trial designed for 80% power with a traditional test gains a 50% chance of stopping early, at roughly the halfway point. If the threshold is not crossed, the trial proceeds to completion and the planned analysis is conducted with no penalty.
This represents a free option: anytime-valid monitoring with no alpha spending and no pre-specified interim looks. The “cost” is that this method alone has lower power than a fixed-sample test. But when layered on top of a properly powered trial, it provides early stopping when effects are larger than anticipated.
@@ 3.3 Trajectory Examples
Figure 1 shows example wealth trajectories from 30 simulated trials under the null hypothesis (both arms 40% event rate). Sample sizes correspond to trials designed for 80% power (n = 712, left) and 90% power (n = 954, right) to detect a 10% ARR. Under the null, wealth fluctuates randomly around 1. Some trajectories temporarily rise toward the threshold but eventually drift back down. The downward drift over time reflects the accumulating “cost” of betting on noise.
Figure 1: Wealth trajectories under the null hypothesis. Left: n = 712 (80% power design). Right: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Dotted gray line: neutral (wealth = 1). Under the null, no trajectory crosses the threshold.